# TreeKV: Smooth Key-Value Cache Compression with Tree Structures

### v2

@@ Abstract
Efficient key-value (KV) cache compression is critical for scaling transformer-based Large Language Models (LLMs) in long sequences and resource-limited settings. Existing methods evict tokens based on their positions or importance scores, but position-based strategies can miss crucial information outside predefined regions, while those relying on global importance scores resulting in strong regional biases, limiting the KV cache’s overall context retention and potentially impairing the performance of LLMs on complex tasks. Our wavelet analysis reveals that as tokens approach the end of sequence, their contributions to generation gradually increase and tends to diverge more from neighboring tokens, indicating a smooth transition with increasing complexity and variability from distant to nearby context. Motivated by this observation, we propose TreeKV, an intuitive, training-free method that employs a tree structure for smooth cache compression. TreeKV maintains a fixed cache size, allowing LLMs to deliver high-quality output even in long text scenarios. Unlike most compression methods, TreeKV is applicable to both the generation and prefilling stages. TreeKV consistently surpasses all baseline models in language modeling tasks on PG19 and OpenWebText2, allowing LLMs trained with short context window to generalize to longer window with a 16x cache reduction. On the Longbench benchmark, TreeKV achieves the best performance with only 6% of the budget at optimal efficiency [^1] We’ll open-source our method upon acceptance..
@@ 4.1 Decoding Stage
10:        $\overline{\mathbf{S}}=\mathbf{S}/\mathbf{C}$. $\triangleright$ Important Scores
@@ 5.3 Long Context Understanding
Table 3 summarizes the performance metrics for all tasks and cache sizes in the Longbench benchmark. Overall, our model achieved an average improvement of 0.74 over H2O and 0.48 over SnapKV, demonstrating its superior capacity to retain and recall relevant information from extended text, it did not surpass the full attention model, which leaves space for us to improve on with minimal resource requirements.
5.4 Ablation Study: What is the Core Factor in TreeKV?

### v3

@@ Abstract
Efficient key-value (KV) cache compression is critical for scaling transformer-based Large Language Models (LLMs) in long sequences and resource-limited settings. Existing methods evict tokens based on their positions or importance, but position-based strategies can miss crucial information outside predefined regions, while those relying on global importance scores resulting in strong regional biases, limiting the KV cache’s overall context retention and potentially impairing the performance of LLMs on complex tasks. Our wavelet analysis reveals that as tokens approach the end of sequence, their contributions to generation gradually increase and tends to diverge more from neighboring tokens, indicating a smooth transition with increasing complexity and variability from distant to nearby context. Motivated by this observation, we propose TreeKV, an intuitive, training-free method that employs a tree structure for smooth cache compression. TreeKV maintains a fixed cache size, allowing LLMs to deliver high-quality output in long text scenarios and is applicable during both the generation and prefilling stages. TreeKV consistently surpasses all baseline models in language modeling tasks on PG19 and OpenWebText2, allowing LLMs trained with short context window to generalize to longer window with a 16x cache reduction. On the Longbench benchmark, TreeKV achieves the best performance with only 6% of the budget at optimal efficiency [^1] https://github.com/ZiweiHe/TreeKV.
@@ 4.1 Decoding Stage
10:        $\overline{\mathbf{S}}=\mathbf{S}/\mathbf{C}$. $\triangleright$ Importance Scores
@@ 5.3 Long Context Understanding
Table 3 summarizes the performance metrics for all tasks and cache sizes in the Longbench benchmark. Overall, our model achieved an average improvement of 0.74 over H2O and 0.48 over SnapKV, demonstrating its superior capacity to retain and recall relevant information from extended text. TreeKV underperforms SnapKV in Multi-Doc QA and Summarization tasks at cache size 2k. Since TreeKV algorithmically prioritizes recent tokens when the cache size is limited, we hypothesize that limiting the tree height to balance early and recent token retention could mitigate this issue, which will explore in future. TreeKV did not surpass the full attention model, which leaves space for us to improve on with minimal resource requirements.
5.4 Ablation Study